# BALF: Budgeted Activation-Aware Low-Rank Factorization for Fine-Tuning-Free Model Compression

### v1

@@ Appendix C Obtaining the Uncentered Whitening Matrices
$R$ being the rank of $\bm{X}$}. Moreover, it is easy to see that $\operatorname{row}({\bm{X}})=\operatorname{range}({\bm{M}})$, as required in our definition (see Section 3). Additionally, the pseudoinverse of $\bm{\Lambda}$} (a diagonal matrix) can be obtained efficiently by inverting each nonzero element of the diagonal, while leaving zeros as is. This avoids additional numerical approximation error when forming $\bm{M}$}. We do not perform any additional steps on the whitening matrices.
@@
We use the linear algebra functions (with CUDA tensors) provided in the PyTorch library. During early experimentation, we used the default solver (cuSOLVER (NVIDIA, 2025)), but we noticed that it sometimes crashed for seemingly well-behaved matrices. Because of that, we switched to the MAGMA (Heroux et al., 2024) backend, which behaved much better. Sporadically, the $\mathsf{eigh}$ routine was still unable to converge (even when the matrix was actually symmetric). For that reason, we added a fallback to the regular $\mathsf{eig}$ routine. We observed that the fallback was needed only for a couple of matrices, and only in the ResNeXt-101 model. In none of our experiments did the fallback also fail (so we do not resort to noise addition in any instance, contrasting with Wang et al. (2025c)).
@@
The naive way of obtaining $\frac{1}{N}{\bm{X}}^{T}{\bm{X}}$ in practice is to use a single batch, and caching activations by passing the batch through the network. However, for some networks with high-dimensional features and/or big batch sizes, the memory requirements might make the process expensive or impossible. A nice property (mentioned in Wang et al. (2025c), but not discussed extensively) is that we can use batched to accumulate the second moments. Suppose we have $M$ batches, then,
@@
Appendix D Rank Allocation Details
@@
Our objective is to select a list $\{P_{l}\}_{l=1}^{L}$ denoting the retained rank in each layer, with the objective of maximizing the global retained energy under some cost constraint. We can model this as a binary linear program in the form of a multiple-choice knapsack problem (for a review, see Szkaliczki (2025)), with decision variables $\{x_{l,t}\}_{l,t=1}^{L,U_{l}}$ where $x_{l,t}\in\{0,1\}$ denotes whether, at layer $l$, $t$ is selected as the maximum retained rank, and $U_{l}$ denotes an upper bound on the rank (in our case, the number of singular values). Furthermore, by establishing a function $C_{l}(P_{l})$ that returns the computational cost (e.g., in FLOPs or number of parameters) of keeping up to rank $P_{l}$ at layer $l$, we can ensure that the complexity of the solution is constrained to some user-set quantity. The binary program can be expressed as:
@@
Recall that we use the total retained energy, which is normalized, as the objective function. Following the insights of Theorem 3 and Theorem 4 (although they are not tight for deeper networks), one might speculate that, with appropriate accounting of each layer’s importance, we could achieve better results. During our initial experiments, we also evaluated unnormalized distortion as a metric in place of energy and explored layer-wise weightings based on statistics such as output variance and entropy. None of these alternatives improved performance over the method presented here; they instead increased complexity and risked additional computational and implementation overhead.
@@
Our allocator makes use of a cost function. We refer the reader to Appendix B for the discussion on the cost of regular and factorized layers, both in terms of FLOPs and in terms of the number of parameters. It is clear that not all $P$ selections result in a fewer number of parameters or FLOPs. Given $P$ and a criterion, we only decompose the layer when we obtain gains in terms of that criterion (e.g., a reduction in FLOPs or parameters); see Algorithm 1. To reflect that in the cost function, we define it as a piecewise linear function, where the first region is monotonically increasing and represents the region where decomposing the layer results in compression, and the second remains constant, denoting that the action of “keeping rank PP” will have the cost of the original layer. Summarizing, we define
$C_{l}(P)=\min\{C_{l}^{\mathrm{orig}},C_{l}^{\mathrm{decomp}}(P)\}.$
@@
D.1 A Scalable Lagrangian Relaxation
@@
Each of the $I$ outer iterations scans all the candidates once, giving time $O(I\sum_{l}U_{l})$ (excluding the initial doubling, which is $O(D\sum_{l}U_{l})$, with $D$ being the number of doublings to find the initial $\lambda_{\max}$, which scales logarithmically with model size). This scales linearly in network size and the number of iterations parameter.
@@
We implement the algorithm in Python, as even for larger models, it has negligible runtime (less than 0.2 seconds in all our experiments, see Appendix H). Moreover, although the energy and cost functions are represented in functional form in the formal analysis, in practice, they are precomputed as vectors for each layer.
Appendix E Bounding the Output Distortion of a Model
@@
There are some interesting instances where the bound reduces to simpler expressions. For example, for fully connected layers, $B_{l}=1$. Moreover, for 1-Lipschitz activations, such as ReLU (Nair & Hinton, 2010) or GELU (Hendrycks & Gimpel, 2016), $A_{l}=1$. For instance, on MLPs with ReLU or GELU activations, the bounds reduce to
@@
Appendix F Experimental Details
@@ F.1 CIFAR-10 Experiments
We train them as follows. We use SGD with momentum 0.9 and weight decay of $10^{-4}$. Training runs for 200 epochs. The learning rate is initialized to 0.1 and reduced by a factor of 0.1 at epochs 100 and 150. We save the final model checkpoint at epoch 200. For data augmentation, we apply 4-pixel zero-padding followed by a random crop to $32\times 32$ pixels, a random horizontal flip with a probability of 0.5, and per-channel normalization.
@@ F.2 ImageNet Experiments
For all the CNNs, we use the implementation provided by the $\mathsf{torchvision}$ library (maintainers & contributors, 2016). We also use their $\mathsf{V1}$ checkpoints for all the models. For the experiments on vision transformers, we use the implementation and weights provided in the $\mathsf{timm}$ library (Wightman, ).
@@
Following standard inference practice, we first resize images to $256\times 256$ and then center-crop to $224\times 224$. We then normalize them as usual. We do this for both the calibration dataset and the test dataset.
F.3 General Details
For all our main experiments, we use 300 iterations for our rank allocator (see Section D.1). For all the compression sweeps in our experiments section, we make a full sweep over the compression targets. For the experiments with our allocator, we select compression ratios (w.r.t. parameter and FLOP counts) between 0.1 and 1.0 in steps of 0.05; additionally, we add 0.975. For the other methods (energy-based methods) we select an appropriate set of hyperparameters for each model so that they cover an appropriate spectrum of compression ratios.
@@ F.3 General Details
F.4 Computational Resources
All the experiments present in the paper were carried out on a laptop with an RTX 2070 and a server with an RTX 4090 (the latter was used mainly for the ImageNet experiments). We note that, for all the models discussed, our framework works and is efficient even on the laptop, but accuracy evaluation is significantly faster on the RTX 4090. Preliminary experimentation that did not make it into the paper was conducted on a variety of hardware configurations, including an A100 40GB and other RTX GPUs.
Appendix G Additional Experimental Results
This section presents some additional experimental results that aim to answer secondary questions. We also include compression-accuracy curves on three models that were not included in the main text due to space constraints.
G.1 How Impactful is the Calibration Dataset Size?
@@
G.2 Robustness under Distribution Shift
Our method relies on the computation of (uncentered) whitening matrices (over the training dataset) in order to estimate the optimal low-rank projections. We rely on the i.i.d. assumption to claim that the statistics of the calibration dataset will be useful to compress a model in a way that transfers to unseen data, and, as shown over the training-test pairs on CIFAR-10 and ImageNet, we obtain good results. In general, the i.i.d. assumption is common (although seldom discussed) in methods that use calibration data.
@@ G.2 Robustness under Distribution Shift
Now, if we evaluate our original (non-compressed) model on the corrupted test set, we can measure the accuracy gap as the difference between the accuracy on the original test set and the corrupted test set: $\Delta^{\text{test-orig}}_{\text{test-corr}}$ (corr stands for “corrupted”). For the sake of simplicity, we assume those are positive (i.e., there are accuracy losses), which is true in general.
@@
We wish to answer the question: How much additional accuracy loss does our compression method induce when evaluated on the corrupted test dataset? Formally, we wish to measure $\Delta^{\text{test-orig}}_{\text{test-comp+corr}}-\Delta^{\text{test-orig}}_{\text{test-corr}}$ (we make the assumption that those are “orthogonal”).
@@
We measure this quantity on the CIFAR-10 dataset. We take our pre-trained ResNet-20 model on CIFAR-10 and compress it across multiple target ratios using BALF (we compress it with different configurations, including FLOPs targets and parameter count targets). We use 1024 CIFAR-10 training images as the calibration set. Additionally, we use the same hyperparameters as in our main experiments. For each resulting compressed model (and for the original model), we then evaluate: (1) its accuracy on the original CIFAR-10 test set, and (2) its accuracy on each corruption and severity level from CIFAR-10-C (Hendrycks & Dietterich, 2019). Over different compression configurations, we report clean accuracy, CIFAR-10-C mean accuracy (over all the corruption configurations), and the gap between them. Finally, we visualize the differences between clean and corrupted accuracies for both models directly and annotate their gap. Results are shown in Figure 6. This allows us to measure the additional loss in robustness introduced by compression on top of the loss already incurred by distribution shift.
@@
G.3 Speedup of Low-Rank Layers in Practice
This section discusses the relationship between our compression results (in terms of FLOPs) and the end-to-end speedup obtained in practice. We use basic PyTorch implementations of low-rank operators. In particular, we substitute each original layer with two sequential layers, as described in Appendix B, without any specialized implementation.
@@ G.3 Speedup of Low-Rank Layers in Practice
We expect that specialized operator implementations, including those found on software for edge devices, will be able to fully exploit the compression in terms of speedup. In fact, Sui et al. (2024), for instance, were able to obtain near-perfect practical speedups from FLOP reductions on ASICs and FPGAs on ResNet-50 with low-rank layers.
@@
We note that this is, in general, a persistent problem of compression methods (see, e.g., Narshana et al. (2023), where, without specialized implementations, the actual speedups are significantly lower than the FLOPs reductions). We leave the development of specialized operators for future work.
G.4 Additional Complexity-Accuracy Figures
@@
Appendix H Runtime and Memory Footprint of BALF
@@
    First, given a calibration dataset, we gather the activation moments. The model runs on the GPU, but the activation moments are offloaded to CPU memory to reduce GPU memory usage and allow compression on lower-end GPUs. Once activation moments are gathered, these can be cached (on disk) for subsequent compression configurations.
@@
    Then, whitening matrices and factors are computed on the GPU. These are cached on disk to maintain low memory usage and to avoid recomputation on future runs.
@@
For a thorough examination of the implementation, we refer the reader to our code  [^4] https://github.com/davidgonmar/BALF. Table 2 presents timings and peak memory of a run of our method on different models on an RTX 2070 laptop. We use 8192 calibration samples (as in our main experiments) divided into batches of 64 samples. The rest of the parameters are also equal to those used in our main experiments. We note that the peak memory usage is usually attained during activation moment computation, and that by using smaller batch sizes, one could further reduce it.
@@
Finally, we reiterate that the most expensive steps (activation moment gathering and factorization/whitening) are cached, allowing one to obtain different models with different compression configurations in a matter of seconds, on average, even on lower-end hardware, after the first run. We also note that our implementation was not particularly optimized (as compression time is not a bottleneck in our evaluations), and that further runtime gains are probably attainable.
Appendix I A Post-Hoc Interpretation via Information Geometry
@@
Let $\widehat{{\bm{\mathsfit{W}}}}$ denote the low-rank approximation of a parameter tensor. They pose that general activation-aware factorization approaches solve an optimization problem equivalent to minimizing the squared output mismatch under an i.i.d. Gaussian noise model with identity covariance. Translating that to our framework, we obtain
@@
Appendix J Missing Proofs
In this section, we provide full proofs for results in the main text that were not covered in other sections of the appendix. Results are restated for the convenience of the reader.
J.1 Proof of Theorem 2

### v2

@@ Appendix B Obtaining the Uncentered Whitening Matrices
$R$ being the rank of $\bm{X}$}. Moreover, it is easy to see that $\operatorname{row}({\bm{X}})=\operatorname{range}({\bm{M}})$, as required in our definition (see Section 3). Additionally, the pseudoinverse of $\bm{\Lambda}$} (a diagonal matrix) can be obtained efficiently by inverting each nonzero element of the diagonal, while leaving zeros as is. This avoids additional numerical approximation errors when forming $\bm{M}$}. We do not perform any additional steps on the whitening matrices.
@@
We use the linear algebra functions (with CUDA tensors) provided in the PyTorch library. During early experimentation, we used the default solver (cuSOLVER (NVIDIA, 2025)), but we noticed that it sometimes crashed for seemingly well-behaved matrices. Because of that, we switched to the MAGMA (Heroux et al., 2024) backend, which behaved much better. Sporadically, the $\mathsf{eigh}$ routine was still unable to converge, even when the matrix was actually symmetric. For that reason, we added a fallback to the regular $\mathsf{eig}$ routine. We observed that the fallback was needed only for a couple of matrices, and only in the ResNeXt-101 model. In none of our experiments did the fallback also fail (so we do not resort to noise addition in any instance, contrasting with Wang et al. (2025c)).
@@
The naive way of obtaining $\frac{1}{N}{\bm{X}}^{T}{\bm{X}}$ in practice is to use a single batch, and then cache activations by passing the batch through the network. However, for some networks with high-dimensional features and/or large batch sizes, the memory requirements might make the process expensive or impossible. A nice property (mentioned in Wang et al. (2025c), but not discussed extensively) is that we can use data batches to accumulate the second moments. Suppose we have $M$ batches, then,
@@
Appendix C Rank Allocation Details
@@
Our objective is to select a list $\{P_{l}\}_{l=1}^{L}$ denoting the retained rank in each layer, with the objective of maximizing the global retained energy under some cost constraint. We can model this as a binary linear program in the form of a multiple-choice knapsack problem (for a review, see Szkaliczki (2025)), with decision variables $\{x_{l,t}\}_{l,t=1}^{L,U_{l}}$, where $x_{l,t}\in\{0,1\}$ denotes whether, at layer $l$, $t$ is selected as the maximum retained rank, and $U_{l}$ denotes an upper bound on the rank (in our case, the number of singular values). Furthermore, by establishing a function $C_{l}(P_{l})$ that returns the computational cost (e.g., in FLOPs or number of parameters) of keeping up to rank $P_{l}$ at layer $l$, we can ensure that the complexity of the solution is constrained to some user-set quantity. The binary program can be expressed as:
@@
Recall that we use the total retained energy, which is normalized, as the objective function. Following the insights of Theorem 3 and Theorem 4 (although they are not tight for deeper networks), one might speculate that, with appropriate accounting of each layer’s importance, we could achieve better results. During our initial experiments, we also evaluated unnormalized distortion as a metric in place of energy and explored layer-wise weightings based on statistics such as output variance and entropy. None of these alternatives improved performance over the method presented here; instead, they increased complexity and posed additional computational overhead and implementation challenges.
@@
Our allocator makes use of a cost function. We refer the reader to Appendix A for the discussion on the cost of regular and factorized layers, both in terms of FLOPs and in terms of the number of parameters. It is clear that not all $P_{l}$ selections result in fewer parameters or FLOPs. Given $P_{l}$ and a criterion, we decompose the layer only when we achieve gains in terms of that criterion (e.g., a reduction in FLOPs or parameters); see Algorithm 1. To reflect this in the cost function, we define it as a piecewise linear function, where the first region is monotonically increasing and represents the region where decomposing the layer results in compression, and the second remains constant, denoting that the action of “keeping rank $P_{l}$” will have the cost of the original layer. Summarizing, we define
$C_{l}(P_{l})=\min\{C_{l}^{\mathrm{orig}},C_{l}^{\mathrm{decomp}}(P_{l})\}.$
@@
C.1 A Scalable Lagrangian Relaxation
@@
Each of the $I$ outer iterations scans all the candidates once, giving time $O(I\sum_{l}U_{l})$. The initial doubling to find $\lambda_{\max}$ adds an additional $O(D\sum_{l}U_{l})$ factor. Thus, the total time complexity is $O((I+D+1)\sum_{l}U_{l})$ (in the case of a nontrivial solution).
@@
We implement the algorithm in Python, as even for larger models, it has a negligible runtime (less than 0.2 seconds in all our experiments, see Appendix G). Moreover, although the energy and cost functions are represented in functional form in the algorithm description, in practice, they are precomputed as vectors for each layer.
Appendix D Bounding the Output Distortion of a Model
@@
There are some interesting instances where the bound reduces to simpler expressions. For example, for fully connected layers, $B_{l}=1$. Moreover, for 1-Lipschitz activations, such as ReLU (Nair & Hinton, 2010) or GELU (Hendrycks & Gimpel, 2023), $A_{l}=1$. For instance, on MLPs with ReLU or GELU activations, the bounds reduce to
@@
Appendix E Experimental Details
@@ E.1 CIFAR-10 Experiments
We use SGD with momentum 0.9, weight decay $10^{-4}$, and a batch size of 128. The learning rate is initialized to 0.1 and reduced by a factor of 0.1 at epochs 100 and 150. We save the final model checkpoint at epoch 200. For data augmentation, we apply 4-pixel zero-padding followed by a random crop to $32\times 32$ pixels, a random horizontal flip with a probability of 0.5, and per-channel normalization. Under these settings, the final test accuracies are 91.93% for ResNet-20 and 93.18% for ResNet-56.
@@ E.2 ImageNet Experiments
For all the CNNs, we use the implementation provided by the $\mathsf{torchvision}$ library (maintainers & contributors, 2016). We also use their $\mathsf{V1}$ checkpoints for all the models. For the experiments on vision transformers, we use the implementation and weights provided in the $\mathsf{timm}$ library (Wightman, 2019).
@@
Following standard inference practice, we first resize images to $256\times 256$ (using bilinear interpolation for all models except for the vision transformers, for which we use bicubic interpolation) and then center-crop them to $224\times 224$. We then normalize them as usual. We do this for both the calibration dataset and the test dataset.
E.3 General Details
For all our main experiments, we use 300 iterations for our rank allocator (see Section C.1). For all the compression sweeps in our experiments section, we conduct a full sweep over the compression targets. For the experiments with our allocator, we select compression ratios (w.r.t. parameter and FLOP counts) between 0.1 and 1.0 in steps of 0.05; additionally, we add 0.975. For the other methods (energy-based methods), we manually select an appropriate set of hyperparameters for each model so that they cover a suitable spectrum of compression ratios.
@@ E.3 General Details
E.4 Computational Resources
All the experiments presented in the paper were conducted on a laptop with an RTX 2070 and a server with an RTX 4090 (the latter was used mainly for the ImageNet experiments). We note that, for all the models discussed, our framework works and is efficient even on the laptop, but accuracy evaluation is significantly faster on the RTX 4090. Preliminary experimentation that did not make it into the paper was conducted on a variety of hardware configurations, including an A100 40GB and other RTX GPUs.
Appendix F Additional Experimental Results
This section presents some additional experimental results that aim to answer secondary questions. We also include compression-accuracy curves for three models that were not included in the main text due to space constraints.
F.1 How Impactful is the Calibration Dataset Size?
@@
F.2 Robustness under Distribution Shift
Our method relies on the computation of (uncentered) whitening matrices (over the training dataset) to estimate the optimal low-rank projections. We rely on the i.i.d. assumption to claim that the statistics of the calibration dataset will be useful for compressing a model in a way that transfers to unseen data. As shown over the training-test pairs on CIFAR-10 and ImageNet, we obtain good results. In general, the i.i.d. assumption is common (although seldom discussed) in methods that use calibration data.
@@ F.2 Robustness under Distribution Shift
Now, if we evaluate our original (non-compressed) model on the corrupted test set, we can measure the accuracy gap as the difference between the accuracy on the original test set and the corrupted test set: $\Delta^{\text{test-orig}}_{\text{test-corr}}$ (corr stands for “corrupted”). For the sake of simplicity, we assume that those are positive (i.e., there are accuracy losses), which is generally true.
@@
We wish to answer the question: How much additional accuracy loss does our compression method induce when evaluated on the corrupted test dataset? Formally, we wish to measure $\Delta^{\text{test-orig}}_{\text{test-comp+corr}}-\Delta^{\text{test-orig}}_{\text{test-corr}}$ (we make the assumption that these are “orthogonal”).
@@
We measure this quantity on the CIFAR-10 dataset. We take our pre-trained ResNet-20 model on CIFAR-10 and compress it across multiple target ratios using BALF (we compress it with different configurations, including FLOPs targets and parameter count targets). We use 1024 CIFAR-10 training images as the calibration set. Additionally, we use the same hyperparameters as in our main experiments. For each resulting compressed model (and for the original model), we then evaluate: (1) its accuracy on the original CIFAR-10 test set, and (2) its accuracy on each corruption and severity level from CIFAR-10-C (Hendrycks & Dietterich, 2019). Over different compression configurations, we report clean accuracy, CIFAR-10-C mean accuracy (over all the corruption configurations), and the gap between them. Finally, we visualize the differences between clean and corrupted accuracies for both models directly and annotate their gaps. Results are shown in Figure 6. This allows us to measure the additional loss in robustness introduced by compression on top of the loss already incurred by distribution shift.
@@
F.3 Speedup of Low-Rank Layers in Practice
This section discusses the relationship between our compression results (in terms of FLOPs) and the end-to-end speedup obtained in practice. We use basic PyTorch implementations of low-rank operators. In particular, we substitute each original layer with two sequential layers, as described in Appendix A, without any specialized implementation.
@@ F.3 Speedup of Low-Rank Layers in Practice
We expect that specialized operator implementations, including those found in software for edge devices, will be able to fully exploit the compression in terms of speedup. In fact, Sui et al. (2024), for instance, were able to obtain near-perfect practical speedups from FLOP reductions on ASICs and FPGAs on ResNet-50 with low-rank layers.
@@
We note that this is, in general, a persistent problem with compression methods (see, e.g., Narshana et al. (2023), where, without specialized implementations, the actual speedups are significantly lower than the FLOPs reductions). We leave the development of specialized operators for future work.
F.4 Additional Complexity-Accuracy Figures
@@
Appendix G Runtime and Memory Footprint of BALF
@@
    First, given a calibration dataset, we gather the activation moments. The model runs on the GPU, but the activation moments are offloaded to CPU memory to reduce GPU memory usage and allow compression on lower-end GPUs. Once activation moments are gathered, they can be cached (on disk) for subsequent compression configurations.
@@
    Then, whitening matrices and factors are computed on the GPU. These are cached on disk to maintain low memory usage and to avoid recomputation in future runs.
@@
For a thorough examination of the implementation, we refer the reader to our code [^4] https://github.com/davidgonmar/BALF. Table 2 presents the runtime breakdown and peak memory usage of our method on different models on the ImageNet dataset, evaluated on an RTX 2070 laptop. We use 8192 calibration samples (as in our main experiments), divided into batches of 64 samples. All other parameters match those used in the main experiments. We note that the peak memory usage typically occurs during the computation of activation moments, and that using smaller batch sizes can further reduce it.
@@
Finally, we reiterate that the most expensive steps (activation moment gathering and factorization/whitening) are cached, allowing one to obtain different models with different compression configurations in a matter of seconds on average, even on lower-end hardware, after the first run. We also note that our implementation was not particularly optimized (as compression time is not a bottleneck in our evaluations), and that further runtime gains are probably attainable.
Appendix H A Post-Hoc Interpretation via Information Geometry
@@
Let $\widehat{{\bm{\mathsfit{W}}}}$ denote the low-rank approximation of a parameter tensor. They posit that general activation-aware factorization approaches solve an optimization problem equivalent to minimizing the squared output mismatch under an i.i.d. Gaussian noise model with identity covariance. Translating that to our framework, we obtain
@@
Appendix I Missing Proofs
In this section, we provide full proofs for the results in the main text that were not covered in other sections of the appendix. Results are restated for the convenience of the reader.
I.1 Proof of Theorem 2